# Unlock the Power of Frozen LLMs in Knowledge Graph Completion

### v2

@@ Abstract
Traditional knowledge graph completion (KGC) methods rely solely on structural information, struggling with the inherent sparsity of knowledge graphs (KGs). Large Language Models (LLMs) learn extensive knowledge from large corpora with powerful context modeling, making them promising for mitigating the limitations of previous methods. Directly fine-tuning LLMs offers great capability but comes at the cost of huge time and memory consumption, while utilizing frozen LLMs yields suboptimal results. In this work, we aim to leverage LLMs for KGC effectively and efficiently. We capture the context-aware hidden states of knowledge triples by employing prompts to stimulate the intermediate layers of LLMs. We then train a data-efficient classifier on these hidden states to harness the inherent capabilities of frozen LLMs in KGC. Additionally, to reduce ambiguity and enrich knowledge representation, we generate detailed entity descriptions through subgraph sampling on KGs. Extensive experiments on standard benchmarks demonstrate the efficiency and effectiveness of our approach. We outperform traditional KGC methods across most datasets and, notably, achieve classification performance comparable to fine-tuned LLMs while enhancing GPU memory efficiency by 188×188$\times$188 × and accelerating training and inference by 13.48×13.48$\times$13.48 ×.
Unlock the Power of Frozen LLMs in Knowledge Graph Completion
Bo Xue, Yi Xu, Yunchong Song, Yiming Pang, Yuyang Ren, Jiaxin Dingthanks:  Jiaxin Ding is the corresponding authors., Luoyi Fu, Xinbing Wang Shanghai Jiao Tong University, Shanghai, China {sappho_x, jiaxinding, yiluofu} @sjtu.edu.cn
1 Introduction
@@
Knowledge graph completion has become a crucial endeavor to improve the application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b). Traditional knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which suffer from the sparseness of KGs (Yao et al., 2019). Leveraging the powerful contextual modeling capabilities and extensive stored world knowledge of LLMs can mitigate this challenge by utilizing the rich semantics within knowledge graphs, thereby enhancing the performance of knowledge graph completion (Wei et al., 2023).
@@
These approaches generally fall into two types: using frozen LLMs with prompting techniques and fine-tuned LLMs specifically for KGC. The first type leverages frozen LLMs with prompt techniques, such as in-context learning. These techniques enable models to perform KGC directly through natural language outputs (Wei et al., 2023; Yao et al., 2023), harnessing the pre-trained knowledge embedded within these models without additional training. While this approach offers computational efficiency, it may not fully capture the task-specific nuances required for optimal KGC performance. The second type involves fine-tuning the LLMs using supervised techniques, including instruction tuning, to adapt models specifically for KGC tasks. This method involves fine-tuning LLMs, such as LLaMA, through the use of training samples (Yao et al., 2023), which can lead to improved performance by aligning the models more closely with the KGC objectives (Zhang et al., 2023; Yao et al., 2023). However, fine-tuning demands significant increases in time and resource expenditure. This dichotomy raises a fundamental question: How can we utilize LLMs to accomplish KGC tasks with both efficacy and efficiency?
To answer this question, we explore the reasons behind the underperformance of frozen LLMs compared to their fine-tuned counterparts. First, while LLMs contain vast amounts of world knowledge encoded within their parameters, accessing the knowledge and decoding it into responses effectively through frozen models can be challenging (Wang et al., 2023). LLMs can produce hallucinations-plausible but incorrect outputs (Zhao et al., 2023)-making them unreliable for precise KGC tasks. Fine-tuning directly modifies the models’ weights through task-specific data enhancing the retrieval of relevant knowledge and improving performance (Zhou et al., 2024). Second, KGs often involve complex relationships and ambiguous entity representations (Li et al., 2024), which can be difficult for a frozen LLM to handle such nuances accurately without specific tuning. Fine-tuned models can be explicitly trained to interpret and disambiguate the complexities enabling better linkage of entities within KGs.
To overcome the limitations of frozen LLMs in KGC tasks, we introduce probing techniques that utilize stimulation prompts, complemented by detailed entity descriptions, as shown in Figure 1. These methods are designed to effectively extract and leverage the latent knowledge embedded within the models’ parameters, and minimize ambiguous entity representations, thereby enhancing its application in KGC without the need for extensive retraining. Specifically, our research concentrates on the core tasks of knowledge graph completion: triple classification and relation prediction.
We stimulate the intermediate layers of LLMs to capture the context-aware hidden states of knowledge triples. Moreover, we improve the precision of knowledge representation by aligning KG entities with the LLMs, reducing the ambiguity of knowledge triples through the generation of entity descriptions via subgraph sampling from KGs. We train a data-efficient classifier on the obtained hidden states suppervised by labels, to fully harness the inherent KGC capabilities of the frozen LLMs.
Figure 1: The overall architecture of our proposed probing method for the triple classification. Constructing positive and negative sample pairs, followed by stimulating the middle layers of the language model to obtain hidden states for completing the KGC task. Entity descriptions are generated by the subgraph entity description generator. The only component that needs training in the entire architecture is the data-efficient classifier used to classify hidden states, which can be MLP, logistic regression, etc.
Our contributions can be summarized as follows.
  • We introduce a novel probing method tailored for KGC tasks, designed to stimulate language models to predict knowledge graph triples directly. We further generate knowledge graph entity descriptions through the use of subgraphs, which are then seamlessly integrated into our probing techniques.
  • Our approach employs LLMs with both effectiveness and efficiency to achieve KGC. By utilizing only a subset of the dataset, our method achieves a relative performance improvement of up to 2.7% compared to fine-tuning the LLM on the entire training set. Additionally, it significantly reduces GPU memory usage by a factor of 188 during the training phase and decreases time consumption by a factor of 13.48, all while maintaining comparable performance. [^1] For a detailed comparison of efficiency, please refer to Table 8 in the Appendix B.3.
  • We conduct extensive experiments on six standard knowledge graph datasets to benchmark our method against multiple baseline approaches. We investigate the versatility of our method with various language models, assessing its efficacy and data efficiency through rigorous experimental validation.
2 Method
Our method is designed based on an analysis of the two primary reasons for the underperformance of frozen LLMs compared to fine-tuned counterparts. First, frozen LLMs may not effectively access the necessary knowledge to generate accurate responses. We demonstrate in Section 2.2 how to achieve this for the frozen LLM using probing techniques. Second, KGs often involve ambiguous entity representations, which can be difficult for LLM to handle accurately. In Section 2.3 we show how to achieve this with the generation of entity descriptions.
2.1 Formulation and Notations
@@
We focus on triple classification and relation prediction. Triple classification involves evaluating the plausibility of a given knowledge triple, to determine whether the triple is true or not and relation prediction predicts the missing relation in a triple.
2.2 Stimulating the Intermediate Layers of Large Language Models
Knowledge graphs consist of numerous triples, and the format of these triples is challenging for the LLM to understand, which may mislead the LLM to retrieve irrelevant knowledge and reveal the wrong responses. To enhance the performance of LLMs in KGC, we leverage the internal judgment mechanisms of the language model (Zou et al., 2023). Our method involves strategically triggering the necessary depth of knowledge embedded within the LLM to elicit the correct responses required for accurate KGC, thereby improving the overall performance of the KGC task.
Specifically, since different layers of the language model have different roles (Azaria and Mitchell, 2023), we determine the most suitable layer for the task by experimenting with the validation set. For the selection of intermediate layers, see Appendix B.5. First, we construct prompt templates to stimulate the model to generate the required classification information in the intermediate layers. In the triple classification, our goal is to determine whether the triples in the knowledge graph are true. Therefore, we use the following prompt template, $PT_{1}$, to create positive and negative sample pairs for training, as shown in Figure 2. Notice that the prompt does not describe positive or negative information for triples.
@@ 2.2 Stimulating the Intermediate Layers of Large Language Models
Then, we use existing triples in the knowledge graph to construct positive samples $S^{+}$, and employ the negative sampling technique (Lv et al., 2022) to obtain negative samples $S^{-}$. Finally, we obtain $N$ training data samples. For each training text, we input it into the language model to obtain the representation of each text in the intermediate layers. We can obtain information about the plausibility of triples from the hidden state corresponding to the token immediately preceding the model prediction, which is the last token of the input text (Zou et al., 2023). Specifically, we capture plausibility information for positive samples $s^{+}\in S^{+}$ by:
$v\left(s^{+}\right)=\{M_{1:i}\left(s^{+}\right)[-1],0<i<L\},$ (1)
where $M$ represents the language model, $M_{1:i}$ represents the first $i$ layers of the model, $[-1]$ refers to the hidden state corresponding to the last token, $L$ denotes the number of layers in the model. The dimension of $M_{1:i}(s^{+})[-1]$ is the embedding dimension of the language model. Negative samples follow the same principle.
At last, we train a classification model using $v(s^{+})$ and $v(s^{-})$ to distinguish between plausible and implausible triples. The goal is to learn a boundary that can separate the hidden states of the intermediate layers of the LLM corresponding to positive samples (which are true) from those corresponding to negative samples (which are false). We can use classification models such as multilayer perceptron (MLP), SVM, etc., for classification training. Taking logistic regression to determine the plausibility of samples as an example:
$P\left(Y=\mathbf{1}|x\right)=\frac{1}{1+e^{-\left(w^{T}x+b\right)}},$ (2)
where x represents $v(s)[l]$, $l$ is the layer index, $w$ and $b$ are weights and bias to be learned.
For relation prediction, we transform the binary classification into multi-class classification. By utilizing the prompt in Figure 7, we stimulate the LLM to generate hidden states that capture the characteristics of different relation types in the KG. Details can be found in the Appendix B.1.
By focusing on extracting classification results directly from the model’s internal processing, we ensure that the responses are both relevant and precise, thereby improving the effectiveness of the KGC task.
Figure 2: Prompt templates to stimulate the LLM for triple classification.
2.3 Subgraph Entity Description Generator
The semantics of entities in the knowledge graph are complex, making it challenging for the language model to interpret based solely on the entity’s name due to inherent ambiguity and polysemy. Therefore, we inform the language model about the meaning of each entity in the input. The meaning of each entity in KG is generated by its one-hop subgraph.
Specifically, we extend the prompt template constructed in Section 2.2, aiming to assist the language model in aligning internal entities with entities in the knowledge graph. For triple classification, since we employ the probing method to extract KGC features from the samples, we need to ensure that the language model can be stimulated to produce effective hidden states that can be used by the classifier to predict the plausibility of triples. Therefore, we construct a new prompt $PT_{2}$, as shown in Figure 2. By introducing the $D(e)$, LLMs are able to generate more precise and distinguishable hidden states, thereby improving the overall classification performance.
@@ 2.3 Subgraph Entity Description Generator
To address incomplete entity descriptions within certain datasets, we have devised a method to generate $D(e)$. We construct $D(e)$ in two ways. The first approach involves a straightforward method where we concatenate all the triples within the one-hop subgraph surrounding the entity $e$, thereby forming $D(e)$ (Zhang et al., 2023). However, this direct concatenation can sometimes lead to ineffective or misleading stimuli for the LLM, as the structure of these subgraph triples may significantly differ from the text seen during the LLM’s pre-training phase. This divergence can adversely affect the model’s response accuracy. To mitigate this issue, our second approach leverages a language model to transform these triples into a format that is more comprehensible to the LLM. By rephrasing the triples from the one-hop subgraph around the entity $e$ into natural language text, we produce a more refined and model-friendly version of $D(e)$:
$D(e)=\operatorname{concat}\{\operatorname{transform}(\{h_{1}\}\{r_{1}\}\{t_{1}\}),$ (3)
$(h_{1},r_{1},t_{1})\in\operatorname{subgraph}(e)\},$
@@
where $\operatorname{subgraph(\cdot)}$ is a function that retrieves the set of all triples in the one-hop subgraph where the entity resides; $\operatorname{transform(\cdot)}$ is a function that converts the triple into sentence; $\operatorname{concat(\cdot)}$ is a function to process triple statements. For the first approach, $\operatorname{concat(\cdot)}$ is a string concatenation function that joins each transformed triple statement with a separator. For the second approach, we need to utilize LLM with the in-context learning technique (Brown et al., 2020) to generate descriptions of entities based on the triples in their subgraph. Please refer to Appendix A.1 for details.
By generating long entity descriptions using the subgraph description generator, we align the internal understanding of entities in the LLM with the entity meanings in the KG, which is similar to the entity linking process in knowledge graphs. In this way, the LLM will better understand the semantics of the entities, and improve the KGC performance.
3 Experiments

### v3

# Bridging LLMs and KGs without Fine-Tuning: Intermediate Probing Meets Subgraph-Aware Entity Descriptions

@@ Abstract.
Traditional knowledge graph completion (KGC) methods rely solely on structural information, struggling with the inherent sparsity of knowledge graphs (KGs). By contrast, Large Language Models (LLMs) encapsulate extensive world knowledge and exhibit powerful context modeling capabilities, making them promising for mitigating the limitations of traditional methods. However, direct fine-tuning of LLMs for KGC, though effective, imposes substantial computational and memory overheads, while utilizing non-fine-tuned LLMs is efficient but yields suboptimal performance. In this work, we propose a novel framework that synergizes the strength of LLMs with robust knowledge representation to enable effective and efficient KGC. We extract the context-aware hidden states of knowledge triples from the intermediate layers of LLMs, thereby capturing rich semantic and relational nuances. These represenation are then utilized to train a data-effecient classifier tailored specifically for KGC tasks. To bridge the semanitic gaps between LLMs and KGs, we employ subgraph sampling on KGs to generate model-friendly entity descriptions, ensuring that the representations are comprehensive and well aligned with the KG structure. Extensive experiments on standard benchmarks validate the efficiency and effectiveness of our approach. We achieve a 47% relative improvement over previous methods based on non-fine-tuned LLMs and, to our knowledge, are the first to achieve classification performance comparable to fine-tuned LLMs while enhancing GPU memory efficiency by 188×188$\times$188 × and accelerating training and inference by 26.11×26.$1$.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Knowledge graph completion has become a crucial endeavor to improve the application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b). Traditional knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which is vulnerable to the sparseness of KGs (Yao et al., 2019). By leveraging the powerful contextual modeling capabilities and extensive stored world knowledge of LLMs, we can effectively mitigate these challenges, and enhance the performance of KGC (Wei et al., 2024).
@@
These approaches generally fall into two types: using non-fine-tuned LLMs with prompting techniques and fine-tuned LLMs specifically for KGC. The first type leverages LLMs with prompt techniques, such as in-context learning. These techniques enable models to perform KGC directly through natural language outputs (Yang et al., 2024; Wei et al., 2024; Yao et al., 2024), harnessing the pre-trained knowledge embedded within these models without additional training. While this approach offers computational efficiency, it may not fully capture the task-specific nuances required for optimal KGC performance. In contrast, the latter type involves fine-tuning the LLMs using supervised techniques to adapt models specifically for KGC tasks. This method involves fine-tuning LLMs, such as LLaMA, through the use of training samples (Yao et al., 2024), which can lead to improved performance by aligning the models more closely with the KGC objectives (Zhang et al., 2024; Yao et al., 2024). However, fine-tuning demands significant increases in time and resource expenditure. This dichotomy between computational efficiency and task-specific performance raises a fundamental question: How can we harness LLMs for effective and efficient KGC without incurring the high costs associated with fine-tuning?
Answering this question necessitates a deeper look at the limitations of LLMs compared to their fine-tuned counterparts. Non-fine-tuned LLMs, while cost-effective, risk generating misleading outputs in the form of hallucinations (Zhao et al., 2023) and underutilizing their latent knowledge (Wang et al., 2023). Moreover, the intricate relational structures and semantic subtleties embedded in KGs compound the challenge, since LLMs are not inherently designed to parse or disambiguate complex entity linkages, without explicitly fine-tuning on such tasks (Li et al., 2024).
In this paper, we propose a novel framework that reconciles these competing demands by fully exploiting the knowledge representation capabilities of LLMs without fine-tuning, as shown in Figure 1. Our method strategically extracts intermediate hidden states from LLMs to capture rich, context-aware representations of knowledge triples, which are then used to train a lightweight, data-efficient classifier specifically tailored for KGC. To bridge the gap between LLMs and KGs, we incorporate an innovative subgraph sampling technique to generate model-friendly entity descriptions, enabling the LLMs to better interpret and disambiguate relational patterns to further enhance knowledge representation. By focusing on these intermediate LLM representations, our framework mitigates hallucination risks while retaining the benefits of a non-fine-tuned paradigm, allowing for both high accuracy and reduced computational costs.
Extensive experiments demonstrate that our framework not only outperforms existing non-fine-tuned LLM-based methods by $47\%$ relative improvement but also achieves classification performance comparable to fine-tuned LLMs. Our approach enhances GPU memory efficiency by a factor of 188×188$\times$188 × and accelerates both training and inference processes by 26.11×26.$1$. Our contributions can be summarized as follows:
Figure 1. The overall architecture of our method for KGC involves categorizing triples based on the type of task, followed by their verbalization. This is combined with entity descriptions, which are generated by the subgraph entity description generator, to produce the stimulation prompt. Finally, the internal representation (hidden states) of the LLM obtained from the stimulation prompt is used to train a classifier, thereby completing the KGC. The only component that needs training in the entire architecture is the data-efficient classifier, making the method easy to train and cheap to deploy.
  • We propose, to the best of our knowledge, the first framework that directly harnesses the intermediate representation space of non-fine-tuned LLMs for KGC tasks. To bridge the semantic gap between the LLM and the KG, we design a subgraph-based entity description generator, ensuring that the model’s internal understanding of entities aligns effectively with KG structures.
  • We provide an in-depth analysis of why fine-tuned LLMs typically achieve superior KGC performance, identifying the core factors responsible for these gains. Building on these insights, we present novel optimization strategies for non-fine-tuned LLMs and are the first to enable them to achieve KGC performance comparable to their fine-tuned counterparts.
  • We perform extensive experiments across standard KGC datasets to validate the effectiveness and efficiency of our proposed method. These substantial improvements highlight the scalability and practical applicability of our method, making it a viable solution for large-scale KGC applications where computational resources and time are critical constraints.
2. Method
Our method is proposed in face of the limitations of non-fine-tuned LLMs on KGC. In Section 2.2, we present how to precisely extract context-aware representation of knowledge from intermediate layers of LLMs. In Section 2.3 we tackle complex relationships in KGs, by the generation of model-friendly entity descriptions through subgraph sampling.
2.1. Formulation and Notations
@@
For triple classification, the task is to assess the plausibility of a given knowledge triple, determining whether the triple is true or not. As for relation prediction and entity prediction, they aim to predict the missing relation and entity in a triple, respectively.
2.2. Extracting Knowledge Representations
Knowledge graphs are composed of numerous triples that encapsulate complex relational and semantic information. For non-fine-tuned LLMs, interpreting these intricate structures poses significant challenges, often leading to the retrieval of irrelevant information and the generation of incorrect or misleading responses.
To enhance the performance of LLMs in KGC, our method leverages the internal judgment mechanisms of LLMs, as shown in (Zou et al., 2023). Specifically, we strategically harness the intermediate layers of LLMs to extract rich, context-aware knowledge representations that are pivotal for accurate KGC.
First, we construct specialized prompt templates designed to stimulate the LLM to generate the necessary classification information within its intermediate layers. For triple classification, where the objective is to determine the plautibility of triples in the KG, we employ a prompt template, $PT_{1}$, that creates positive and negative sample pairs for training, as illustrated in Figure 2. Notably, the prompt avoids explicitly describing the positivity or negativity of triples.
@@ 2.2. Extracting Knowledge Representations
Thereafter, we use existing triples in the knowledge graph to construct positive samples $S^{+}$, and employ the negative sampling technique (Lv et al., 2022) to obtain negative samples $S^{-}$. This results in a dataset obtaining $N$ training samples. Each training sample is input into the LLM to obtain representations from its intermediate layers.
Specifically, we capture plausibility information and extract hidden states for positive samples $s^{+}\in S^{+}$ as follows:
(1) $Rep_{j}^{i}\left(s^{+}\right)=\{M_{1:i}\left(s^{+}\right)\left[j\right],0\textless i\textless L,0\textless j\textless N\},$
where $M$ represents the language model, $M_{1:i}$ represents the first $i$ layers of the model, $[j]$ refers to the hidden state corresponding to the $j$th token, $L$ and $N$ represent the number of layers in the model and the number of tokens resulting from tokenizing $s^{+}$. Consistent with findings in (Zou et al., 2023), we utilize the representation of the last token in each input text, as it is more conducive to information extraction due to the autoregressive nature of causal language models. Negative samples are processed similarly to ensure consistency in representation extraction.
The extracted representations $Rep_{j}^{i}(s^{+})$ and $Rep_{j}^{i}(s^{-})$ serve as input features for training a lightweight classifier to distinguish between plausible and implausible triples by learning a decision boundary that effectively separates the hidden states of positive samples from those of negative samples. Although the classifiers employed—such as multilayer perceptrons (MLPs) and logistic regression models—are inherently simple, their effectiveness is significantly amplified by the quality and richness of the intermediate layer representations, thereby enhancing the ability to make accurate predictions without the computational overhead associated with fine-tuning the entire LLM.
For relation prediction, we extend our binary classification framework to a multi-class setting, where the classifier predicts the specific relationship types within the KG. Utilizing the $PT_{3}$ shown in Figure 2, we stimulate the LLM to generate hidden states that capture the characteristics of different relation types. This approach ensures that the classifier can effectively learn to differentiate between multiple relationship categories from the model’s internal processing, enhancing the accuracy of KGC.
Figure 2. Prompt templates to stimulate the LLM.
2.3. Subgraph Entity Description Generator
The semantics of entities in a KG can be complex and ambiguous, posing significant challenges for LLMs to interpret accurately based solely on entity names. To overcome these challenges and enhance the semantic comprehension of entities by LLMs without resorting to computationally intensive fine-tuning, we introduce the subgraph entity description generator, which enriches entity representations by providing additional contextual descriptions derived from their local subgraph structures. Specifically, we extend the prompt template constructed in Section 2.2. We focus on aligning the language model’s internal entity understanding with those in the knowledge graph. Therefore, we construct new prompts $PT_{2}$ and $PT_{4}$, as shown in Figure 2. By introducing the entity description $D(e)$, LLMs are able to generate more precise and distinguishable hidden states, thereby improving the overall classification performance.
@@ 2.3. Subgraph Entity Description Generator
Most KGs contain entity descriptions that can be directly used as $D(e)$. However, to address incomplete entity descriptions within certain KG datasets, we propose a method to generate $D(e)$. Unlike the approach in (Yang et al., 2024), which generates the entity description using only a single triple, our method constructs a more precise $D(e)$ by incorporating subgraph information in two distinct ways. For the first approach, we concatenate all the triples within the one-hop subgraph around the entity $e$, thereby forming $D(e)$ (Zhang et al., 2024). However, providing information in the form of explicit rules may be less effective for LLM to retrieve and leverage knowledge (Bian et al., 2024), as the structure of these subgraph triples may significantly differ from the text seen during the pre-training phase. Therefore, our second approach leverages a language model to transform these triples into a narrative format that is more comprehensible to the LLM. By rephrasing the triples from the one-hop subgraph around the entity $e$ into natural language text, we produce a more refined and model-friendly version of $D(e)$ that is more effective for LLMs to retrieve and leverage knowledge. Specifically:
(2) $D(e)=\operatorname{Rephrase}(\{\operatorname{Verbalize}(\{h_{i}\}\{r_{i}\}\{t_{i}\}),$
$(h_{i},r_{i},t_{i})\in\operatorname{Subgraph}(e)\}),$
@@
where $\operatorname{Subgraph(\cdot)}$ is a function that automatically retrieves a specified number of triples in the one-hop subgraph where the entity resides; $\operatorname{Verbalize(\cdot)}$ is a function that transform the triple into its textual string, in this work, we use the linear verbalization (Baek et al., 2023); For the first approach, $\operatorname{Rephrase(\cdot)}$ is a string concatenation function that joins each verbalized triple statement with a separator. For the second approach, we employ a LLM with the in-context learning technique, then ask it to generate model-friendly descriptions of entities based on the verbalized triples in their subgraph.
By leveraging the subgraph entity description generator, we align the LLM’s internal entity understanding with those in the KG, akin to entity linking process. This alignment not only mitigates ambiguities but also enhances LLM’s semantic comprehension, ultimately boosting KGC performance.
Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
TransE S-F ALL 0.815 0.759 0.697 0.884 0.845
DistMult S-F ALL 0.862 0.871 0.587 0.851 0.864
ComplEx S-F ALL 0.857 0.862 0.657 0.841 0.871
RotatE S-F ALL 0.819 0.847 0.685 0.882 0.877
KG-BERT L-MLP ALL 0.899 0.935 0.560 0.916 0.773
KGT5 N-L ALL 0.663 0.728 - - -
LLAMA-7B N-L 0 0.699 0.661 0.573 0.458 0.658
LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
KG-LLAMA-7B N-L ALL 0.892 0.955 0.748 0.921 0.858
KG-LLAMA-SAMPLED N-L 3k/10k 0.727 0.682 0.614 0.574 0.834
LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.934 0.862
LLAMA-DES2 (Tri) H-MLP 3k/10k 0.842 0.755 0.671 0.779 0.782
LLAMA-DES2 (GPT) H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
LLAMA-DES-H H-MLP ALL 0.920 - 0.753 0.943 0.863
Table 1. The main experiment results (accuracy) of triple classification. In the ‘Pred’ column, $S\mbox{-}F$ uses score function, $L\mbox{-}MLP$ uses MLP on last layer hidden states, $N\mbox{-}L$ uses natural language, and $H\mbox{-}MLP$ uses MLP on intermediate layer hidden states to predict. ‘Samples’ indicates training sample counts, with $3k/10k$ indicating $10k$ samples for the UMLS dataset and $3k$ for others. LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimulus generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimulus generation. Top three results per dataset are bolded.
3. Experiments